OpenStudy (anonymous):

Please help will give medal If tan x° = 11 divided by r and cos x° = r divided by s, what is the value of sin x°? A. sin x° = s divided by 11 B. sin x° = 11 divided by s C. sin x° = 11r D. sin x° = 11s

OpenStudy (hoblos):

you have to use the identity \[\tan x = \frac{ \sin x }{ \cos x }\]

OpenStudy (hoblos):

replace tanx and cosx by their values what do you get ?

OpenStudy (anonymous):

so it would be B then right?

OpenStudy (hoblos):

correct answer

OpenStudy (anonymous):

thank you!

OpenStudy (hoblos):

glad to help :)